OpenStudy (anonymous):

A rectangular box is twice as long as it is wide. The height of the box is 3 feet less than the width. If the box is x feet wide, what polynomial represents its volume in cubic feet?

OpenStudy (anonymous):

Let width of rectangular box be x feet. Then the length is 2x feet. Height of the box given = x - 3 feet Volume of a cuboid = length*width*height Volume of a cuboid = x*2x*(x-3) cubic. feet Volume of the cuboid: \[2x ^{2}*(x-3) = 2x ^{3} - 6x ^{2}\] cubic feet. which is the final answer.